OpenStudy (anonymous):

Which statement best explains the relationship between current, voltage, and resistance? If we increase the amount of voltage applied, and do not change the resistance, this will result in a decrease in current. If we decrease the current applied, and do not change the resistance, we increase the voltage. If we increase the amount of voltage applied, and do not change the resistance, we will also increase the current. If we decrease the amount of current, this will not affect the amount of voltage, only the amount of resistance.

OpenStudy (compphysgeek):

The best statement is Ohm's Law: \[ V = R * I\] Voltage is the product of Resistance and Current, or \[I = \frac{V}{R}\] if we increase the voltage the current increases as resistance is constant.\[R = \frac{V}{I}\] if voltage is constant and current drops then resistance must be increased. Ohm's Law says it all

OpenStudy (geekfromthefutur):

So basically its If we decrease the amount of current, this will not affect the amount of voltage, only the amount of resistance.